# Personalized Item Representations in Federated Multimodal Recommendation

### v2

@@ 4.1. Encoding Multimodal Information with FM
Eq. 3 has two main advantages: 1) It transforms the visual or textual modality $\mathbf{E}_{m}$ into a format that the recommendation model can process; 2) Compared to transmitting raw modality, the encoded feature embeddings require less communication overhead, thus improving communication efficiency. Moreover, FedMR also builds an id embedding $\mathbf{D}$ to capture the structure information of items.
@@ 4.2. Mixing Feature Fusion Module
Considering the diverse interests and behavior patterns of users, FedMR incorporates a Mixing Feature Fusion Module (MFFM) that adopts multiple modality fusion strategies $\{g_{j}\}$ from $\mathcal{G}$ to fuse the multimodal feature embeddings $\mathbf{V}\text{, }\mathbf{C}$, and the id embedding $\mathbf{D}$ :
(4) $\mathbf{F}_{j}=g_{j}(\mathbf{V},\mathbf{C},\mathbf{D};\gamma_{j}),$
where $\gamma_{j}$ is parameters of the fusion strategy $g_{j}$, and $\mathbf{F}_{j}$ is the feature embedding obtained after applying the $j$-th fusion strategy. To capture user diversity by enabling each strategy to focus on specific user groups based on their interaction histories $\mathbf{r}_{u}$, FedMR introduces a user-specific router $\phi_{u}$, driven by the parameters $\varphi_{u}$:
(5) $\mathbf{w}_{u}=\phi_{u}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|};\mathbf{r}_{u},\varphi_{u}).$
The router $\mathbf{r}_{u}$ dynamically generates a weight vector $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ based on the feature embeddings of items the user has interacted with, obtained through different fusion strategies, and assigns weights to each fusion strategy, i.e., the weight of the strategy $g_{j}$ is $w_{u,j}$. Then, we have the final fused features $\bar{\mathbf{F}}^{(u)}$ as follows:
@@
According to Eq. (6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item representations $\bar{\mathbf{F}}^{(u)}$, thereby enhancing the system’s ability to capture fine-grained user preferences.
@@ 4.3. Objective of FedMR
By replacing $\mathbf{D}$ in Eq. (2) with $\bar{\mathbf{F}}$ obtained from Eq. (6) and then substituting into Eq. (1), we have the objective of FedMR as follows:
(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}}^{(u)};\theta_{u}),r_{u,i}).$
Eq. (7) defines the objective of FedMR and is compatible with most existing ID-based FedRec methods. By replacing the original reconstruction error calculation function and the values of $\{\alpha_{u}\}$ in these methods with Eq. (7), they can integrate the multimodal information of items into personalized recommendations while protecting user privacy. This leads to a more comprehensive understanding of item features and enhances the model’s grasp of user preferences.
4.4. Framework
Fig. 1 illustrates the overall framework of FedMR, which leverages a FM to process multimodal information for enhancing personalized recommendations. On the server side, FedMR uses item cover images as visual modality data and titles as textual modality data. These inputs are processed by the FM to generate fixed visual embeddings $\mathbf{V}$ and text embeddings $\mathbf{C}$. Since both the FM and the input data are fixed, these item feature embeddings do not need to be updated during training. Additionally, the server maintains an ID embedding $\mathbf{D}$ to capture the structural information of the items.
Upon receiving the features $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM. Within this module, a learnable user-specific router $\phi_{u}$ dynamically assigns weights to the fused embeddings for further weighted fusion. The fusion strategies include a Sum strategy, a MLP (Taud and Mas, 2018) strategy, and a Gate strategy. The personalized item representations $\bar{\mathbf{F}}^{(u)}$ are then fed into a prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$, and optimize the model parameters by minimizing reconstruction loss in Eq. (7).
During the entire process, only the ID embedding $\mathbf{D}$ and the gradients of fusion strategies in the MFFM are sent back to the server for aggregation and updates, hereby protecting user privacy while integrating shared information across clients. The core of FedMR lies in dynamically adjusting weights in the MFFM based on the user’s local interactions $\mathbf{r}_{u}$, thereby enhancing the diversity and personalization of the RecSys while ensuring privacy protection. As a result, FedMR is compatible with existing ID-based FedRec methods, allowing multimodal item information to be integrated into these methods. Furthermore, FedMR can be extended by incorporating more modalities or adjusting fusion strategies as needed.
@@ 4.4. Framework
3:     $S_{t}$leftarrow randomly select $n_{s}$ from $n$ clients;
4:     for all client index $u\in S_{t}$ do
5:        $\hat{\mathbf{r}}_{u}$, $\mathbf{D}^{(u)}$, $\{\nabla^{(u)}_{j}\gamma\}$leftarrow ClientUpdate($\varphi$, $\theta$);
6:     end for
7:     $\mathbf{D}\leftarrow\sum_{u\in S_{t}}\alpha_{u}\mathbf{D}^{(u)}$;
8:     $\gamma_{j}\leftarrow\gamma_{j}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla^{(u)}_{j}\gamma$, $j=1,2,\cdots,|\mathcal{G}|$;
9:  end for
10:  return: $\hat{\mathbf{R}}=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n}]^{T}$
ClientUpdate:
1:  $\nabla^{(u)}_{j}{\gamma}\leftarrow 0$, $j=1,2,\cdots|\mathcal{G}|$;
2:  for $a=1,2,\dots,A$ do
3:     Compute partial gradients $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$ and $\{\nabla^{(u)}_{\gamma_{j}}\}$ by differentiating Eq. (7);
4:     Update $\theta_{u}\leftarrow\theta_{u}-\eta\nabla_{\theta_{u}}$;
5:     Update $\varphi_{u}\leftarrow\varphi_{u}-\eta\nabla_{\varphi_{u}}$;
6:     Update $\mathbf{D}\leftarrow\mathbf{D}-\eta\nabla_{\mathbf{D}}$;
7:     Update $\gamma_{j}\leftarrow\gamma_{j}-\eta\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
8:     Accumulate $\nabla^{(u)}_{j}{\gamma}\leftarrow\nabla^{(u)}_{j}{\gamma}+\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
@@
11:  return: $\hat{\mathbf{r}}_{u}$, $\mathbf{D}$, $\{\nabla^{(u)}_{j}{\gamma}\}$
4.5. Algorithm
To optimize the FedMR objective function defined in Eq. (7), we propose an alternating optimization algorithm for training FedMR. The detailed steps of the algorithm are presented in Alg. 1.
The FedMR framework consists of two main components: the GlobalProcedure and the ClientUpdate. In the GlobalProcedure, the server first uses a FM $h(\cdot)$ to encode the multimodal data, generating the corresponding embeddings $\mathbf{V}$ and $\mathbf{C}$. The server then randomly selects a subset of clients to participate in each training round to reduce the risk of privacy leakage caused by clients participating in consecutive rounds (Li et al., 2023a). For each selected client $u\in S_{t}$, the client update process is executed to obtain the local prediction result $\hat{\mathbf{r}}_{u}$, the local updated id embeddings $\mathbf{D}^{(u)}$, and the local accumulated gradient $\{\nabla^{(u)}_{j}\gamma\}$. The server aggregates the results returned by the clients in a weighted manner, and updates the global shared embedding $\mathbf{D}$ and fusion strategy parameters $\{\gamma_{j}\}$.
During the ClientUpdate, each client calculates the gradients of the model parameters, such as $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$, and $\nabla^{(u)}_{\gamma_{j}}$, based on the FedMR objective function in Eq. (7) and updates these parameters using gradient descent (Amari, 1993). Additionally, each client accumulates the local gradient $\nabla^{(u)}_{\gamma_{j}}$ into $\nabla^{(u)}_{j}\gamma$ for the global update of the fusion strategy parameters. By incorporating a FM $h(\cdot)$ to handle multimodal data $\mathcal{M}$, FedMR achieves efficient personalized multimodal recommendation while preserving user privacy.
@@ 4.6. Complexity Analysis
Since FedMR is compatible with various FedRec methods, we focus on the time and space complexity introduced by incorporating multimodal data of items in this work. Assuming $n$ users, $m$ items and $q$ experts, FedMR , where the multimodal information of the items is encoded on the server side into feature embeddings $\mathbf{V}\in\mathbb{R}^{m\times d}$ and $\mathbf{C}\in\mathbb{R}^{m\times d}$ using a foundation model, with $d\ll\{n,m\}$. Since these embeddings remain fixed during training, they contribute a space complexity of $O(md)$ while having negligible impact on time complexity. FedMR also maintains an ID embedding $\mathbf{D}\in\mathbb{R}^{m\times d}$ to capture structural information of the items, with a space complexity of $O(md)$. During each communication round, the server aggregates and updates local data from the clients, requiring a time complexity of $O(\alpha_{u}md(q+1))$, while each client performs the Mixing Feature Fusion Module to fuse features, resulting in a total client-side space complexity of $O(md+K)$ and a time complexity of $O(P)$, where $K$ and $P$ represents the size and cost time of the component parameters, such as Router, MLP, Gate, etc, respectively. Therefore, FedMR requires a total space complexity of $O(n(md+K)+3md)$ and a total time complexity of $O(nmd(q+1)+nP)$.
Table 1. The statistical information of the datasets used in our work includes the following: #Interactions represents the number of observed interactions. #Users denotes the number of users. #Items indicates the number of items.

### v3

@@ 4.2. Server-Side Multimodal Encoding
This server-side encoding yields communication-efficient, high-quality multimodal representations. Alongside these content embeddings $\mathbf{V}$ and $\mathbf{C}$, FedMR maintains the standard item ID embedding matrix $\mathbf{D}$ to capture collaborative filtering signals (Liu et al., 2023). The three representations $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ then used to form the input for the subsequent fusion module of FedMR.
@@ 4.3. Mixing Feature Fusion Module
FedMR introduce a novel Mixing Feature Fusion Module (MFFM), designed to dynamically fuse the server-provided global representations with client-specific context. MFFM operates in two stages:
Server-Side Multi-Strategy Fusion.
To capture diverse feature interactions, the server employs a predefined set of $|\mathcal{G}|$ distinct fusion strategies $\{g_{j}\}_{j=1}^{|\mathcal{G}|}$. Each strategy $g_{j}$, parameterized by $\gamma_{j}$, combines the multimodal embeddings $\mathbf{V}$ and $\mathbf{C}$, and the ID embedding $\mathbf{D}$ to generate a specific fused representation $\mathbf{F}_{j}$:
(4) $\mathbf{F}_{j}=g_{j}(\mathbf{V},\mathbf{C},\mathbf{D};\gamma_{j}).$
Using multiple strategies produces a diverse set of feature representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$, offering various perspectives on item characteristics. These representations are then transmitted to the clients.
Client-Side Personalized Refinement.
To address user heterogeneity, each client $u$ employs a personalized refinement mechanism. A user-specific router module $\phi_{u}$, parameterized by local parameters $\varphi_{u}$, dynamically computes importance weights $\mathbf{w}_{u}$ for the received server representations based on the client’s interaction history $\mathbf{r}_{u}$:
(5) $\mathbf{w}_{u}=\phi_{u}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|};\mathbf{r}_{u},\varphi_{u}),$
where $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ is typically normalized. Each weight $w_{u,j}$ signifies the relevance of the $j$-th feature perspective $\mathbf{F}_{j}$ for user $u$. The final personalized item representation $\bar{\mathbf{F}}^{(u)}$ is obtained by adaptively combining the server representations using these learned weights:
@@
This adaptive weighting, learned locally based on $\mathbf{r}_{u}$, allows FedMR to tailor global representations to individual user preferences and local data characteristics, directly mitigating the impact of non-IID data and enhancing personalization with privacy-preserving.
@@ 4.4. Objective of FedMR
By substituting the representation $\bar{\mathbf{F}}_{i}^{(u)}$ corresponding to item $i$ for the standard ID embedding $\mathbf{D}_{i}$ within the prediction function $f$ in Eq. (2), the overall objective in Eq. (1) becomes the joint optimization of global and local parameters to obtain the predicted scores $\hat{r}_{u,i}$:
(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}}_{i}^{(u)};\theta_{u}),r_{u,i}).$
Eq. (7) maintains compatibility with existing FedRec frameworks using prediction functions like $f$, allowing them to benefit from richer, personalized, multimodal item representations while preserving privacy. This leads to a more comprehensive understanding of item features and enhances the model’s grasp of user preferences.
Figure 1. The framework of FedMR. It comprises two components: (1) Server-Side Multi-Strategy Fusion, which fuses multimodal features on server, and (2) Client-Side Personalized Refinement, which dynamically adjusts representations based on user finteractions for personalized recommendations on each client while preserving user’s data privacy.
@@
3:    $S_{t}\leftarrow$ randomly select $n_{s}$ from $n$ clients;
4:    Compute $\mathbf{F}_{j}$ according to Eq. (4), $j=1,2,\cdots|\mathcal{G}|$;
5:    for all client index $u\in S_{t}$ do;
6:         $\hat{\mathbf{r}}_{u}$, $\nabla_{\mathbf{D}}^{(u)}$, $\{\nabla^{(u)}_{j}\gamma\}\leftarrow\text{ClientUpdate}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|})$;
7:    end for
8:    $\mathbf{D}\leftarrow\mathbf{D}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla_{\mathbf{D}}^{(u)}$;
9:    $\gamma_{j}\leftarrow\gamma_{j}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla^{(u)}_{j}\gamma$, $j=1,2,\cdots,|\mathcal{G}|$;
10:end for
11:return: $\hat{\mathbf{R}}=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n}]^{T}$
PersonalizedRefinement:
1:$\nabla_{\mathbf{D}}^{(u)}\leftarrow 0$; $\triangleright$ Initialize gradient accumulators
[…]
3:for $a=1,2,\dots,A$ do
4:    Compute partial gradients $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$ and $\{\nabla_{\gamma_{j}}\}$ by differentiating Eq. (7);
5:    Update $\theta_{u}\leftarrow\theta_{u}-\eta\nabla_{\theta_{u}}$triangleright Update local prediction parameters;
6:    Update $\varphi_{u}\leftarrow\varphi_{u}-\eta\nabla_{\varphi_{u}}$triangleright Update local router parameters;
7:    Accumulate $\nabla_{\mathbf{D}}^{(u)}\leftarrow\nabla_{\mathbf{D}}^{(u)}+\nabla_{\mathbf{D}}$triangleright Accumulate gradients;
8:    Accumulate $\nabla^{(u)}_{j}{\gamma}\leftarrow\nabla^{(u)}_{j}{\gamma}+\nabla_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
@@
11:return: $\hat{\mathbf{r}}_{u}$, $\nabla_{\mathbf{D}}^{(u)}$, $\{\nabla^{(u)}_{j}{\gamma}\}$
4.5. Unified Framework and Algorithm
The FedMR framework, illustrated in Fig. 1, integrates multimodal information into FedRecs while preserving privacy through an alternating optimization procedure, as shown in Alg. 1). In each communication round $t$, the server first computes the multi-strategy fused representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$ in Eq. (4) using current global parameters ($\mathbf{D},\{\gamma_{j}\}$) and transmits them to a randomly selected client subset $S_{t}$. Each client $u\in S_{t}$ then utilizes its personalized router $\phi_{u}$ in Eq. (5) to compute user-specific weights $\mathbf{w}_{u}$ based on its local history $\mathbf{r}_{u}$, generating the personalized item representation $\bar{\mathbf{F}}^{(u)}$ in Eq. (6). Using $\bar{\mathbf{F}}^{(u)}$, the client performs local updates on its parameters $\theta_{u}$ and $\varphi_{u}$ by minimizing the objective in Eq. (7) on its private data. Concurrently, it computes gradients w.r.t. the global parameters, $\nabla_{\mathbf{D}}^{(u)}$ and $\{\nabla^{(u)}_{\gamma_{j}}\}_{j=1}^{|\mathcal{G}|}$. These gradients are accumulated and sent back to the server. Finally, the server aggregates the received gradients from all clients in $S_{t}$, typically weighted by $\alpha_{u}$, and updates the global parameters $\mathbf{D}$ and $\{\gamma_{j}\}$. This iterative exchange enables collaborative learning. Crucially, the client-side personalized refinement directly addresses data heterogeneity, while privacy is maintained as only gradients are communicated. FedMR’s design thus flexibly integrates multimodal data, enables fine-grained personalization, maintains compatibility with ID-based methods, and ensures efficiency and accuracy within privacy constraints.
4.6. Convergence Analysis
Here, we provide a theoretical analysis of the FedMR under assumptions for non-convex federated optimization. For simplicity, we denote our learnable parameters as $\mathbf{w}=(\Theta_{g},\{\Theta_{l,u}\}_{u\in\mathcal{U}})$, where $\Theta_{g}=\{\mathbf{D},\{\gamma_{j}\}\}$ represents the set of global parameters, and $\Theta_{l,u}=\{\theta_{u},\varphi_{u}\}$ denotes the local parameters for client $u$, capturing user-specific prediction and router parameters. The global objective is:
(8) $F(\mathbf{w})\;=\;\sum_{u\in\mathcal{U}}\alpha_{u}\,F_{u}(\mathbf{w}),$
where each $F_{u}(\mathbf{w})$ is the local objective associated with user $u$, and $\alpha_{u}>0$ with $\sum\alpha_{u}=1$. We aim to find $\mathbf{w}$ that minimizes $F(\mathbf{w})$.
To establish the convergence guarantees, we rely on the following assumptions commonly used in the analysis of FL algorithms:
Assumption 1 (Smoothness).
Each local objective $F_{u}(\mathbf{w})$ is $L$-smooth w.r.t. $\mathbf{w}$. Hence, the global objective $F$ is also $L$-smooth:
(9) $\|\nabla F_{u}(\mathbf{w})-\nabla F_{u}(\mathbf{w}^{\prime})\|\;\leq\;L\,\|\mathbf{w}-\mathbf{w}^{\prime}\|.$
Assumption 2 (Bounded Gradient Variance).
For the stochastic gradient $g_{l,u}(\mathbf{w};\xi)$ of $F_{u}$, we have $\mathbb{E}_{\xi}[g_{l,u}]=\nabla_{\Theta_{l,u}}F_{u}(\mathbf{w})$ and $\mathbb{E}_{\xi}[\|g_{l,u}-\nabla_{\Theta_{l,u}}F_{u}(\mathbf{w})\|^{2}]\leq\sigma_{l}^{2}$. Furthermore, there exists $G>0$ such that $\|\nabla F_{u}(\mathbf{w})\|\leq G$ in expectation across clients.
Assumption 3 (Bounded Heterogeneity).
The variance of all local gradients across clients is bounded as follows:
$\mathbb{E}_{u}\bigl{[}\|\nabla F_{u}(\mathbf{w})-\nabla F(\mathbf{w})\|^{2}\bigr{]}\;\leq\;\sigma^{2}.$
These assumptions ensure that each $F_{u}$ is reasonably well-behaved and client heterogeneity is bounded. Next, we formalize local drift lemmas and provide the main theorem on FedMR’s convergence.
Let $A$ be the local number of SGD steps per communication round. During each round $t$, a subset of clients $\{u\in S_{t}\}$ updates their local parameters $\Theta_{l,u}$ using a local learning rate $\eta$, while the global parameters $\Theta_{g}$ remain fixed until aggregation. We can show:
Lemma 0 (Local Drift Bound).
Under Assumptions 1 and 2, the expected deviation in local parameters after $A$ updates is bounded by a term proportional to $A\eta^{2}(G^{2}+\sigma_{l}^{2})$.
Lemma 1 induces a gradient bias when uploading local updates to the server. Combining the drift bound with smoothness yields:
Lemma 0 (Gradient Bias Bound).
Let $\bar{g}_{g}^{t}$ denote the averaged gradient for global parameters $\Theta_{g}$ after $A$ local updates on the sampled clients, and $B_{t}=\mathbb{E}[\bar{g}_{g}^{t}|\mathbf{w}^{t}]-\nabla_{\Theta_{g}}F(\mathbf{w}^{t})$ be the gradient bias due to local updates relative to the global parameters $\mathbf{w}^{t}$ at the start of the round. Under Assumption 1 and Lemma 1, We then have:
(10) $\mathbb{E}[\|B_{t}\|^{2}]\leq L^{2}C_{1}A\eta^{2}(G^{2}+\sigma_{l}^{2}).$
We now can show FedMR converges to a stationary point of $F(\mathbf{w})$ in a manner consistent with typical non-convex FL rates:
Theorem 3 (Convergence of FedMR).
Suppose Assumptions 13 hold and let $\eta=c_{\eta}/\sqrt{T}$ for a suitable constant $c_{\eta}>0$. Then for $T$ communication rounds, the FedMR algorithm satisfies:
(11) $\small\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\bigl{[}\|\nabla F(\mathbf{w}^{t})\|^{2}\bigr{]}\;\leq\;O\Bigl{(}\frac{1}{\sqrt{T}}\Bigr{)}\;+\;O\Bigl{(}\frac{AG^{2}}{T}\Bigr{)}\;+\;O\Bigl{(}\frac{\sigma^{2}+G^{2}}{\sqrt{T}\,n_{s}}\Bigr{)}.$
Hence, FedMR achieves a convergence rate of $O(T^{-1/2})$ to a stationary point under non-convex settings, matching the standard rates in federated optimization. The above theorem indicates that FedMR’s multimodal feature fusion and personalization do not degrade the asymptotic rate, and it remains comparable to baseline FL methods. The result also quantifies how larger local steps $A$, stronger data heterogeneity $\sigma^{2}$, or fewer sampled clients $n_{s}$ can slow practical convergence. Nonetheless, for moderate $A$, stable $\sigma^{2}$, and typical $n_{s}$, FedMR effectively converges and benefits from the rich multimodal item representations under the federated setup.
@@ 4.7. Complexity Analysis
In FedMR, the server initially encodes multimodal features into fixed embeddings $\mathbf{V},\mathbf{C}\in\mathbb{R}^{|\mathcal{I}|\times d}$ using an FM $h(\cdot)$, contributing primarily $O(|\mathcal{I}|d)$ space complexity each, as this is often pre-computed. The server also stores the learnable ID embedding $\mathbf{D}\in\mathbb{R}^{|\mathcal{I}|\times d}$, adding $O(|\mathcal{I}|d)$ space. In each round, the server computes $|\mathcal{G}|$ fused representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$ and aggregates gradients, and the dominant per-round time complexity on the server is $O(|\mathcal{G}||\mathcal{I}|d)$.
On the client side, storing local parameters $K$ and temporarily holding the $|\mathcal{G}|$ received representations requires approximately $O(|\mathcal{I}|d+K)$ space. The client’s per-iteration time complexity for local updates, involving the prediction function $f$ and the MFFM router $\phi_{u}$, is denoted as $O(P)$. Therefore, the total system space complexity is roughly $O(|\mathcal{U}|(|\mathcal{I}|d+K)+|\mathcal{I}|d)$, and the total time complexity per round involves server computation plus aggregated client computation, scaling approximately as $O(|\mathcal{G}||\mathcal{I}|d+n_{s}AP)$ where $n_{s}=|S_{t}|$ and $A$ is the number of local steps. This analysis indicates manageable overhead for incorporating FedMR to achieve personalized and privacy-preserving multimodal recommendation.
4.8. Privacy Preservation
FedMR ensures user privacy primarily through the inherent data localization of the federated learning paradigm, where raw interaction data $\mathbf{r}_{u}$ remains on the client’s device. Communication is limited to model gradients or updates. The MFFM is designed to preserve this standard: server-side computations involve only global representations, while the client-specific router $\phi_{u}$ and its parameters $\varphi_{u}$ remain entirely local. The gradients $\nabla\mathbf{D}^{(u)}$ and $\{\nabla^{(u)}_{j}\gamma\}$ shared with the server are analogous to those shared in standard FedRecs, introducing no additional direct data leakage channels and aligning with current privacy practices (Li et al., 2023a; Zhang et al., 2023c; Li et al., 2024c; Feng et al., 2024; Wu et al., 2024). Privacy is further enhanced by standard techniques like random client sampling in each round (Chai et al., 2020; Li et al., 2023a). Moreover, FedMR is fully compatible with advanced privacy-enhancing technologies such as differential privacy (DP) (Abadi et al., 2016), which can be applied by adding noise to gradients before transmission, and other secure aggregation protocols. Our ablation studies provide empirical evidence that DP can be incorporated without drastically compromising the performance.
5. Experiments
5.1. Datasets
We evaluate FedMR’s performance on four public datasets from the NineRec benchmark [^1] Publicly available at: https://github.com/westlake-repl/NineRec (Zhang et al., 2024a): KU, Bili_Food (Food), Bili_Dance (Dance), and Bili_Movie (Movie). Detailed statistics are provided in Table 1. The KU dataset contains short video interactions from the KuaiShou platform. The Food, Dance, and Movie datasets comprise video interactions from distinct categories on the BiliBili platform. These datasets feature diverse content and user interaction patterns but share the challenge of extreme data sparsity (over 95%). We chose these four datasets from the NineRec benchmark for comprehensive evaluation. They represent diverse real-world scenarios, originating from two distinct platforms and covering varied content domains at different scales. This diversity, coupled with their inherent high sparsity, provides a robust testbed for evaluating FedMR’s adaptability and performance across challenging, realistic conditions. Following standard practice, we filter out users with fewer than 5 interactions to mitigate cold-start issues. For data preprocessing, missing item images are imputed using the average visual features of existing items, and missing item titles are replaced with the placeholder text ”The title is missing.”
Table 1. The statistical information of the datasets used in our work includes the following: #Interactions represents the number of observed interactions. #Users denotes the number of users. #Items indicates the number of items. Sparsity is the percentage of #Interactions out of the total possible ratings.